### v1

@@ 5 Open Questions and Further Research
We have introduced the concept of unbent collections, but we leave many questions open. For example, can we efficiently decide, for a given 4-plane graph $G$, whether $\operatorname{un}(G)=2$ or $\operatorname{un}(G)=3$? Does every plane 4-graph $G$ admit an unbent collection of size $\operatorname{un}(G)$ that minimizes $\operatorname{tbn}(G)$? We have shown how to efficiently compute the unbent bend number of triconnected cubic graphs, but there are other natural classes to explore. For example, cubic graphs that are not triconnected, bipartite planar graphs, and series-parallel 4-graphs are good candidates as they admit unbent collections of size 2.
@@ Acknowledgments.
Appendix 0.A Omitted Proofs of Section 2
See 1
Proof
@@ Appendix 0.A Omitted Proofs of Section 2
When we recursively continue decomposing each $G_{C}$ with respect to a cut $X$, we ignore the edges of $X$ when searching for a 2-edge cut in all subsequent iterations. Since every proper induced subgraph $G^{\prime}$ of $G$ with at least $12|V(G^{\prime})|-$ 1 edges is connected via at most a 2-edge cut to the rest of $G$, the base case of this recursion consists of graphs that themselves have at least $12n-$ 1 edges (ignoring the degree-1 vertices) or graphs that have no such high-density subgraphs. Because the latter graphs always admit an unbent collection of size at most 2 by Proposition 2, we can ignore them.
Since, for every remaining component $C$, the corresponding graph $G_{C}$ may still have the degree-1 vertices and, therefore, does not necessarily have at least ${2|V(G_{C})|-1}$ edges overall, we attach to each degree-1 vertex a gadget $H$ (see Fig. 2) of constant size with $12|V(H)|-$ 1 edges and $\operatorname{un}(G+e)\leq 2$, where $e$ is the edge incident to the degree-1 vertex. This ensures that every resulting $n$-vertex 4-graph has at least $12n-$ 1 edges, which together compose $\mathcal{G}$.
@@
See 4
Proof
By Theorem 2.1 we know that $\operatorname{un}(G)\leq 3$. Suppose $\operatorname{un}(G)\leq 2$, then by Proposition 3 $G$ admits a balanced 2-edge-coloring.
@@
Appendix 0.B Omitted Proofs of Section 3
0.B.1 NP-Completeness of Determining the Unbent Bend Number
@@
We start by proving that the problem LevelPlanarMonotone2in4Sat is NP-complete. In 2in4Sat, the input is a Boolean formula in conjective normal form where every clause consists of four literals, and the task is to decide whether an assignment of variables to Boolean values exists such that in every clause exactly two literals are true. In LevelPlanarMonotone2in4Sat, additionally, the variable–clause incidence graph is planar and admits a drawing where the variable vertices are placed on a set of parallel lines, called levels, and the clause vertices are placed between the levels such that no edge crosses a level. The levels divide the plane into rectangular strips. The clauses that are represented in the first, third, etc. strip contain only negated literals while the clauses represented in the second, fourth, etc. strip contain only unnegated literals.
See 2
Proof
The membership in NP is obvious. To show the NP-hardness, we reduce from the NP-hard problem PlanarPositive2in4Sat [10], which is 2in4Sat with the additional property that the incidence graph is planar and all literals are unnegated. As a first step, we reduce to LevelPlanarPositive2in4Sat, where we require a leveling as in LevelPlanarMonotone2in4Sat but all literals are unnegated. Consider a planar straight-line drawing of the incidence graph, which exists and can be computed in polynomial time [7, 14]. Rotate the drawing slightly such that no pair of a variable vertex and a clause vertex lies on the same y-coordinate. Take the set of horizontal lines going through the variable vertices as levels. [^1] In the resulting leveling, every variable vertex may lie on its own level. There may be edges that cross levels. We now show how to reduce the number of edge–level crossings by one in each iteration. For an illustration, see Fig. 7.
@@
See 3.1
Proof
Containment in NP is clear. We show NP-hardness by reduction from LevelPlanarMonotone2in4Sat. Let $\Phi$ be the given $\mathsf{2in4Sat}$ formula, let $m$ be the number of clauses in $\Phi$, let $n$ be the number of variables in $\Phi$, let $H$ be the given embedded (level-planar) variable–clause incidence graph of $\Phi$, and let $\ell$ be the number of levels of $H$. We now construct a plane 4-graph $G$ that follows the structure of the incidence graph $H$. The general structure repeats at every level. Figure 8 illustrates the construction for one level of variable vertices; the non-negated clauses are above that level and the negated clauses are below.
Figure 8: NP-hardness reduction: A layer of variable gadgets and the neighboring clause gadgets. (This figure is a copy of Fig. 5 for the reader’s convenience.)
@@
In $G$, there is a face for each variable vertex (variable gadget; see Fig. 10), for each clause vertex (clause gadget; see Fig. 11), and for each set of edges that are incident and above (below) the same vertex (tunnel gadget). The gadgets are attached to each other according to $H$. The gadgets are designed such that they need to receive specific (non-negative) flow values in Tamassia’s flow model. They are surrounded by a set of faces with demand 0 called walls, which make it costly for flow to enter the gadgets unless it is through some door. (All walls have “thickness” $2n$, that is, each unit of flow traversing a wall bends at least $2n$ edges.) We construct each face with a specific demand. Most faces have demand 0 (so they can be drawn without bends as they are), the clause gadgets have demand 2, and an extra face in each level has demand 1 (called level gadget, shown on the right in Fig. 8). These demands sum up to $C=2m+\ell$. The external face is the only face with a negative demand (namely, $-C$). We will guarantee that there is an unbent collection (with two drawings) and at most $k=2nC$ bends if and only if $\Phi$ is a yes-instance. The key property is that the flow needs to avoid walls, which are, in terms of bends, too expensive to traverse, and hence the flow must traverse the variable and tunnel gadgets to enter the clause gadgets and the level gadgets. Since every variable gadget has one door to enter and two exit doors to leave the gadget (the latter being single edges), in the one drawing it must take the one exit door and in the other drawing it must take the other exit door. This choice corresponds to an assignment of Boolean values to variables.
@@
A $k$-valve gadget is a plane subgraph of $G$ that is a path of $k+2$ vertices where each of the $k$ inner vertices gets a leaf attached. Along the path, all leaves are attached on the same side. The path separates two faces – we call the one containing the degree-1 vertices source face and the other face target face. Since each inner vertex of a $k$-valve gadget has degree 3, it has one free angle, which can be assigned either to the source or to the target face. In the base state, the target face receives all free angles of the inner vertices (see Figs. 9(a) and 9(c)), whereas in the $t$-bent state, the source face receives $t$ free angles (and the target face receives $k-t$ free angles) from the inner vertices where $t\in[k]$ (see Figs. 9(b) and 9(d)). If $k$ is clear from the context or irrelevant, we sometimes call a $k$-valve gadget just valve gadget. For the sake of presentation, we assume that all valve gadgets are by default in the base state.
@@
There are $n$ variable gadgets – one per variable. For an illustration of a variable gadget together with the attachment to the tunnel gadgets above and below, see Fig. 10. A variable gadget has a central face (light blue background) that is bounded by walls except for an $(2m+1)$-valve gadget (shown in dark blue on the left) and two exit doors, which are two single edges (shown in red incident to the light blue area). All faces have demand 0, however, to allow flow to connect the external face with the clause gadgets and the level gadgets without going through walls (which is too expensive in terms of bends), some flow must enter, and the same flow value must leave each variable gadget. The arrangement of variable gadgets (see Fig. 8) requires that flow from the external face or from the variable gadget to the left enters at the $(2m+1)$-valve gadget and leaves at the exit doors. Note that at most one exit door can be bent per drawing and, hence, all flow traversing the variable gadget leaves trough the same exit door in a drawing. On the other hand, the $(2m+1)$-valve gadget can be used in all drawings of the unbent collection as long as at most $t\leq 2m+1$ units of flow are involved (then it is in the $t$-bent state instead of the base state and does not create bends on edges).
@@
There are $2n$ tunnel gadgets – two per variable. For an illustration of tunnel gadgets (shown as slim tunnels with light brown background), see Fig. 8. For the connection to a variable gadget, see Fig. 10, and, for the connection to a clause gadget, see Fig. 11. Each tunnel gadget connects an exit door of a variable gadget with all negated or with all unnegated clause gadgets such that this variable occurs in the corresponding clause. Like all of our gadgets, it is bounded by walls. Flow that enters the tunnel gadget can leave it immediately via a $(2m+1)$-valve gadget to a face leading to the next variable gadget or to a level gadget (described below; see the large yellow faces in Fig. 8). This way flow from the external face can reach all variable gadgets. The path to the clause gadget is intermitted by a set of edges (depicted as the red horizontal line segments in Figs. 8 and 10 that are not the exit doors of variable gadgets). The number of these edges depends on which variable gadget is affected: a tunnel gadget of the $i$-th leftmost variable gadget on a level has $n-i$ such edges. In this way we ensure that every unit of flow reaching a clause gadget from the external face causes at least $n$ bends: $i$ bends for traversing $i$ variable gadgets and $n-i$ bends for traversing a tunnel gadget of the $i$-th variable gadget of some level.
@@
There are $\ell$ level gadgets – one per level. For an illustration of a level gadget, see the rightmost (purple) face with a “1” in Fig. 8. It is bounded by walls except for two doors with $n-s(j)$ edges each (shown in red on the left side of the level gadget) where $s(j)$ is the number of variables assigned to level $j$. These doors connect the level gadget with the face following the $s(j)$-th variable gadget. Since it has a demand of 1, any flow coming from the external face effects at least $n$ bends – by going first through all variable gadgets of the level and then entering the level gadget via one of its two doors. The purpose of the level gadget is to bend an exit door of each variable gadget of this level if traversing walls is avoided. Otherwise, if, say, the rightmost variable gadget does not forward any flow to its adjacent clause gadgets in the first drawing, then it would not bend any exit door in the first drawing and could bend both exit doors in the second drawing.
@@
There are $m$ clause gadgets – one per clause. For an illustration of a clause gadget together with the attachment to the tunnel gadgets, see Fig. 11. A clause gadget is bounded by walls but has four doors connecting it to four tunnel gadgets via a 1-valve gadget. These four tunnel gadgets lead to the four corresponding variable gadgets of the clause. A clause gadget has a demand of 2, which can be satisfied without creating any bends if and only if exactly two of the 1-valve gadgets are in the 1-bent state.
@@
0.B.2 Star Arboricity of Bounded-Degree Graphs
See 5
Proof
@@ 0.B.2 Star Arboricity of Bounded-Degree Graphs
Appendix 0.C Ommitted Proofs of Section 4
@@
Let $G$ be a cubic triconnected plane graph with $n$ vertices. Observe that computing an orthogonal representation of a graph with the minimum number of bends is equivalent to subdividing the edges of the graph the minimum number of times so that the graph admits an orthogonal representation without bends. Hence, our strategy will be to subdivide the edges of $G$ the minimum number of times so that the subdivided graph admits an orthogonal representation with no bend. We shall often call dummy a subdivision vertex and we shall often use the expression “place $h$ dummies along and edge” to mean that the edges has been subdivided $h$ times. Also, in the following $f_{\operatorname{ext}}$ denotes the external face of $G$.
0.C.1 Bad Cycles and Demanding Cycles
@@
A bad cycle is a $k$-legged cycle where $k\leq 3$ and such that less than $4-k$ dummies are placed along its edges. If a cycle is not a bad cycle, it is a good cycle. For example: a 3-legged cycle is bad if no dummy is placed along an edge of it; the external face, which is a 00-legged cycle, is a bad cycle if less than $4$ dummies are placed along its edges. The following lemma gives a sufficient and necessary condition for a graph to be admit an orthogonal drawing without bends and also states that, in this case, there exists a liner-time algorithm computing an orthogonal representation of such graph with no bend.
See 4
Let $e=uv$ be an edge of the external face $f_{\operatorname{ext}}$ of $G$ and let $C$ be the cycle forming the external boundary of $G\setminus e$. Observe that, after placing one or more dummies along $e$, cycle $C$ becomes a (possibly degenerate) 2-legged cycle. See for example Fig. 12. The following lemma gives a sufficient condition for such 2-legged cycles and for degenerate cycles to not exist.
@@
See for example Fig. 12, where the conditions of Lemma 6 hold and $C_{1}$ and $C_{2}$ are good cycles, as they both have two dummies along their edges. Thanks to Lemma 6, we can focus our attention mainly on 3-legged cycles and observe that the other bad cycles (degenerate or 2-legged) are satisfied by the way we place dummies.
Let $C$ be a 3-legged cycle of $G$. We say that the three paths of $C$ connecting the vertices of the cycle incident to the legs are the contour paths of $C$. A face incident to a leg of a 3-legged cycle is called leg face. Note that a 3-legged cycle has exactly three leg faces. Cycles $C_{1},...,C_{8}$ of Fig. 13(a) are all 3-legged cycles. Let $G_{C}$ be the subgraph of $G$ induced by $C$ and its interior vertices and let $n_{C}$ the number of vertices of $G_{C}$. We say that a cycle $C$ contains another cycle $C^{\prime}$ if $C$ is a subgraph of $G_{C}$. The genealogical tree $T(C)$ of a 3-legged cycle $C$ is a tree such that: The root of the tree is $C$; the children of $C$ are the 3-legged cycles of $G_{C}$ such there is no other cycle $C^{\prime}\subset G_{C}$ containing them. Fig. 13(a) shows $T(C_{1})$. Two 3-legged cycles intersect if they have an edge in common, but there is no containment relationship between them.
@@
Cycle $C^{\prime}$ is demanding if its red-green coloring obeys Rule 1 and non-demanding otherwise. For an example of application of the coloring rule, see Fig. 13(a).
Let $C$ be a cycle incident to the external face $f_{\operatorname{ext}}$, i.e., sharing at least an edge with $f_{\operatorname{ext}}$. The twin of $C$, denoted as $C^{t}$, is the 3-legged cycle having the same legs of $G$. Notice that also $C^{t}$ is a cycle incident to $f_{\operatorname{ext}}$. Let $I(G)$ be the set of demanding cycles of $G$ that are intersecting.
Lemma 7 ([5])
@@
See for example Fig. 13(b), which depicts a orthogonal drawing with no bends and with minimum dummies of the graph in Fig. 13(a), where dummies are represented with a cross. If dummies are removed with bends the drawing has the minimum number of bends.
0.C.2 A Better Lower Bound for tbn(G)
In this section we refine a lower bound for $tbn(G)$ proving Lemma 5. In Section 0.C.3 we prove that such lower bound is tight. Before proving Lemma 5 we prove and observe intermediate results.
Observe that, for each unbent collection $\mathcal{C}$ of $G$, $|\mathcal{C}|\geq 2$, as the external face of $G$ is a 0-legged cycle, and by Lemma 4, we have to place at least four dummies in $G$. We introduce the following corollary of Lemma 8.
Corollary 1

### v2

@@ 5 Open Questions and Further Research
We have introduced the concept of unbent collections. We propose here several open questions. Can we efficiently decide, for a given 4-plane graph $G$, whether $\operatorname{un}(G)=2$ or $\operatorname{un}(G)=3$? Does every plane 4-graph $G$ admit an unbent collection of size $\operatorname{un}(G)$ that minimizes $\operatorname{tbn}(G)$? We have shown how to efficiently compute the unbent bend number of plane triconnected cubic graphs, but there are other natural classes to explore. For example, plane cubic graphs that are not triconnected, bipartite planar graphs, and series-parallel 4-graphs are good candidates as they admit unbent collections of size 2.
Lastly, we discuss a subtlety of Definition 1.1. Throughout the paper, we have required the drawings in the unbent collection to have the same embedding as that of the given plane 4-graph. This requirement stems from the NP-hardness of the bend minimization problem in the variable-embedding setting [8]. However, we suggest for future exploration a variant in which one does not have to adhere to the same embedding in all the drawings in the collection; we call a collection allowing for variable embedding mutable. Additionally, we propose a variant between the two definitions where we still allow only one fixed embedding which is however not given in the input; we call a corresponding collection uniform. Observe that the mutable unbent bend number is different from the unbent bend number even considering our introductory example of $K_{4}$. Indeed, the mutable collection can contain two copies of the left drawing in Figure 1(b). Hence, its mutable unbent bend number is at most 10 (as opposed to $\operatorname{tbn}(K_{4})=12$).
@@ Acknowledgments.
Appendix A Omitted Proofs of Section 2
See 2.7
Proof A.1.
@@ Appendix A Omitted Proofs of Section 2
When we recursively continue decomposing each $G_{C}$ with respect to a cut $X$, we ignore the edges of $X$ when searching for a 2-edge cut in all subsequent iterations. Since every proper induced subgraph $G^{\prime}$ of $G$ with at least $12|V(G^{\prime})|-$ 1 edges is connected via at most a 2-edge cut to the rest of $G$, the base case of this recursion consists of graphs that themselves have at least $12n-$ 1 edges (ignoring the degree-1 vertices) or graphs that have no such high-density subgraphs. Because the latter graphs always admit an unbent collection of size at most 2 by Proposition 2.5, we can ignore them.
Since, for every remaining component $C$, the corresponding graph $G_{C}$ may still have the degree-1 vertices and, therefore, does not necessarily have at least ${2|V(G_{C})|-1}$ edges overall, we attach to each degree-1 vertex a gadget $H$ (see Figure 2) of constant size with $12|V(H)|-$ 1 edges and $\operatorname{un}(G+e)\leq 2$, where $e$ is the edge incident to the degree-1 vertex. This ensures that every resulting $n$-vertex 4-graph has at least $12n-$ 1 edges, which together compose $\mathcal{G}$.
@@
See 2.10
Proof A.3.
By Theorem 2.3 we know that $\operatorname{un}(G)\leq 3$. Suppose $\operatorname{un}(G)\leq 2$, then by Proposition 2.9 $G$ admits a balanced 2-edge-coloring.
@@
Appendix B Omitted Proofs of Section 3
B.1 NP-Completeness of Determining the Unbent Bend Number
@@
We start by proving that the problem LevelPlanarMonotone2in4Sat is NP-complete. In 2in4Sat, the input is a Boolean formula in conjective normal form where every clause consists of four literals, and the task is to decide whether an assignment of variables to Boolean values exists such that in every clause exactly two literals are true. In the problem LevelPlanarMonotone2in4Sat, additionally, the variable–clause incidence graph is planar and admits a drawing where the variable vertices are placed on a set of parallel lines, called levels, and the clause vertices are placed between the levels such that no edge crosses a level. The levels divide the plane into rectangular strips. The clauses that are represented in the first, third, etc. strip contain only negated literals while the clauses represented in the second, fourth, etc. strip contain only unnegated literals.
See 3.1
Proof B.1.
The membership in NP is obvious. To show the NP-hardness, we reduce from the NP-hard problem PlanarPositive2in4Sat [10], which is 2in4Sat with the additional property that the incidence graph is planar and all literals are unnegated. As a first step, we reduce to LevelPlanarPositive2in4Sat, where we require a leveling as in LevelPlanarMonotone2in4Sat but all literals are unnegated. Consider a planar straight-line drawing of the incidence graph, which exists and can be computed in polynomial time [5, 14]. Rotate the drawing slightly such that no pair of a variable vertex and a clause vertex lies on the same y-coordinate. Take the set of horizontal lines going through the variable vertices as levels. [^1] In the resulting leveling, every variable vertex may lie on its own level. There may be edges that cross levels. We now show how to reduce the number of edge–level crossings by one in each iteration. For an illustration, see Figure 7.
@@
See 3.2
Proof B.2.
Containment in NP is clear. We show NP-hardness by reduction from the problem LevelPlanarMonotone2in4Sat. Let $\Phi$ be the given 2in4Sat formula, let $m$ be the number of clauses in $\Phi$, let $n$ be the number of variables in $\Phi$, let $H$ be the given embedded (level-planar) variable–clause incidence graph of $\Phi$, and let $\ell$ be the number of levels of $H$. We now construct a plane 4-graph $G$ that follows the structure of the incidence graph $H$. The general structure repeats at every level. Figure 8 illustrates the construction for one level of variable vertices; the non-negated clauses are above that level and the negated clauses are below.
Figure 8: NP-hardness reduction: A layer of variable gadgets and the neighboring clause gadgets. Numbers inside faces (including the external face) indicate the demands of the faces. This figure is a copy of Figure 5 for the reader’s convenience.
@@
In $G$, there is a face for each variable vertex (variable gadget; see Figure 10), for each clause vertex (clause gadget; see Figure 11), and for each set of edges that are incident and above (below) the same vertex (tunnel gadget). The gadgets are attached to each other according to $H$. The gadgets are designed such that they need to receive specific (non-negative) flow values in Tamassia’s flow model. They are surrounded by a set of faces with demand 0 called walls, which make it costly for flow to enter the gadgets unless it is through some door. (All walls have “thickness” $2n$, that is, each unit of flow traversing a wall bends at least $2n$ edges.) We construct each face with a specific demand. Most faces have demand 0 (so they can be drawn without bends as they are), the clause gadgets have demand 2, and an extra face in each level has demand 1 (called level gadget, shown on the right in Figure 8). These demands sum up to $C=2m+\ell$. The external face is the only face with a negative demand (namely, $-C$). In other words, the external face has supply $C$ and no other face supplies flow. We will guarantee that there is an unbent collection (with two drawings) and at most $k=2nC$ bends if and only if $\Phi$ is a yes-instance. The key property is that the flow needs to avoid walls, which are, in terms of bends, too expensive to traverse, and hence the flow must traverse the variable and tunnel gadgets to enter the clause gadgets and the level gadgets. Since every variable gadget has one door to enter and two exit doors to leave the gadget (the latter being single edges), in the one drawing it must take the one exit door and in the other drawing it must take the other exit door. This choice corresponds to an assignment of Boolean values to variables.
@@
A $k$-valve gadget is a plane subgraph of $G$ that is a path of $k+2$ vertices where each of the $k$ inner vertices gets a leaf attached. Along the path, all leaves are attached on the same side. The path separates two faces – we call the one containing the degree-1 vertices source face and the other face target face. Since each inner vertex of a $k$-valve gadget has degree 3, it has one free angle, which can be assigned either to the source or to the target face. In the base state, the target face receives all free angles of the inner vertices (see Figures 9(a) and 9(c)), whereas in the $t$-bent state, the source face receives $t$ free angles (and the target face receives $k-t$ free angles) from the inner vertices where $t\in[k]$ (see Figures 9(b) and 9(d)). If $k$ is clear from the context or irrelevant, we sometimes call a $k$-valve gadget just valve gadget. For the sake of presentation, we assume that all valve gadgets are by default in the base state.
@@
There are $n$ variable gadgets – one per variable. For an illustration of a variable gadget together with the attachment to the tunnel gadgets above and below, see Figure 10. A variable gadget has a central face (light blue background) that is bounded by walls except for an $(2m+1)$-valve gadget (shown in dark blue on the left) and two exit doors, which are two single edges (shown in red incident to the light blue area). All faces have demand 0, however, to allow flow to connect the external face with the clause gadgets and the level gadgets without going through walls (which is too expensive in terms of bends), some flow must enter, and the same flow value must leave each variable gadget. The arrangement of variable gadgets (see Figure 8) requires that flow from the external face or from the variable gadget to the left enters at the $(2m+1)$-valve gadget and leaves at the exit doors. Note that at most one exit door can be bent per drawing and, hence, all flow traversing the variable gadget leaves trough the same exit door in a drawing. On the other hand, the $(2m+1)$-valve gadget can be used in all drawings of the unbent collection as long as at most $t\leq 2m+1$ units of flow are involved (then it is in the $t$-bent state instead of the base state and does not create bends on edges).
@@
There are $2n$ tunnel gadgets – two per variable. For an illustration of tunnel gadgets (shown as slim tunnels with orange background), see Figure 8. For the connection to a variable gadget, see Figure 10, and, for the connection to a clause gadget, see Figure 11. Each tunnel gadget connects an exit door of a variable gadget with all negated or with all unnegated clause gadgets such that this variable occurs in the corresponding clause. Like all of our gadgets, it is bounded by walls. Flow that enters the tunnel gadget can leave it immediately via a $(2m+1)$-valve gadget to a face leading to the next variable gadget or to a level gadget (described below; see the large yellow faces in Figure 8). This way flow from the external face can reach all variable gadgets. The path to the clause gadget is intermitted by a set of edges (depicted as the red horizontal line segments in Figures 8 and 10 that are not the exit doors of variable gadgets). The number of these edges depends on which variable gadget is affected: a tunnel gadget of the $i$-th leftmost variable gadget on a level has $n-i$ such edges. In this way we ensure that every unit of flow reaching a clause gadget from the external face causes at least $n$ bends: $i$ bends for traversing $i$ variable gadgets and $n-i$ bends for traversing a tunnel gadget of the $i$-th variable gadget of some level.
@@
There are $\ell$ level gadgets – one per level. For an illustration of a level gadget, see the rightmost (purple) face with a “1” in Figure 8. It is bounded by walls except for two doors with $n-s(j)$ edges each (shown in red on the left side of the level gadget) where $s(j)$ is the number of variables assigned to level $j$. These doors connect the level gadget with the face following the $s(j)$-th variable gadget. Since it has a demand of 1, any flow coming from the external face effects at least $n$ bends – by going first through all variable gadgets of the level and then entering the level gadget via one of its two doors. The purpose of the level gadget is to bend an exit door of each variable gadget of this level if traversing walls is avoided. Otherwise, if, say, the rightmost variable gadget does not forward any flow to its adjacent clause gadgets in the first drawing, then it would not bend any exit door in the first drawing and could bend both exit doors in the second drawing.
@@
There are $m$ clause gadgets – one per clause. For an illustration of a clause gadget together with the attachment to the tunnel gadgets, see Figure 11. A clause gadget is bounded by walls but has four doors connecting it to four tunnel gadgets via a 1-valve gadget. These four tunnel gadgets lead to the four corresponding variable gadgets of the clause. A clause gadget has a demand of 2, which can be satisfied without creating any bends if and only if exactly two of the 1-valve gadgets are in the 1-bent state.
@@
B.2 Star Arboricity of Bounded-Degree Graphs
See 3.4
Proof B.3.
@@ B.2 Star Arboricity of Bounded-Degree Graphs
Appendix C Ommitted Proofs of Section 4
@@
Let $G$ be a plane triconnected cubic graph with $n$ vertices. Observe that computing an orthogonal representation of a graph with the minimum number of bends is equivalent to subdividing the edges of the graph the minimum number of times so that the graph admits an orthogonal representation without bends. Hence, our strategy will be to subdivide the edges of $G$ the minimum number of times so that the subdivided graph admits an orthogonal representation with no bend. We shall often call dummy a subdivision vertex and we shall often use the expression “place $h$ dummies along and edge” to mean that the edges has been subdivided $h$ times. Also, in the following $f_{\operatorname{ext}}$ denotes the external face of $G$.
C.1 Bad Cycles and Demanding Cycles
@@
A bad cycle is a $k$-legged cycle where $k\leq 3$ and such that less than $4-k$ dummies are placed along its edges. If a cycle is not a bad cycle, it is a good cycle. For example: a 3-legged cycle is bad if no dummy is placed along an edge of it; the external face, which is a 00-legged cycle, is a bad cycle if less than $4$ dummies are placed along its edges. The following lemma gives a sufficient and necessary condition for a graph to admit an orthogonal drawing without bends and also states that, in this case, there exists a liner-time algorithm computing an orthogonal representation of such graph with no bend.
See 4.1
Let $e=uv$ be an edge of the external face $f_{\operatorname{ext}}$ of $G$ and let $C$ be the cycle forming the external boundary of $G\setminus e$. Observe that, after placing one or more dummies along $e$, cycle $C$ becomes a (possibly degenerate) 2-legged cycle. See for example Figure 12. The following lemma gives a sufficient condition for such 2-legged cycles and for degenerate cycles to not exist.
@@
See for example Figure 12, where the conditions of Lemma C.1 hold and $C_{1}$ and $C_{2}$ are good cycles, as they both have two dummies along their edges. Thanks to Lemma C.1, we can focus our attention mainly on 3-legged cycles and observe that the other bad cycles (degenerate or 2-legged) are satisfied by the way we place dummies.
Let $C$ be a 3-legged cycle of $G$. We say that the three paths of $C$ connecting the vertices of the cycle incident to the legs are the contour paths of $C$. A face incident to a leg of a 3-legged cycle is called leg face. Note that a 3-legged cycle has exactly three leg faces. Cycles $C_{1},...,C_{8}$ of Figure 13(a) are all 3-legged cycles. Let $G_{C}$ be the subgraph of $G$ induced by $C$ and its interior vertices and let $n_{C}$ the number of vertices of $G_{C}$. We say that a cycle $C$ contains another cycle $C^{\prime}$ if $C$ is a subgraph of $G_{C}$. The genealogical tree $T(C)$ of a 3-legged cycle $C$ is a tree such that: The root of the tree is $C$; the children of $C$ are the 3-legged cycles of $G_{C}$ such there is no other cycle $C^{\prime}\subset G_{C}$ containing them. Figure 13(a) shows $T(C_{1})$. Two 3-legged cycles intersect if they have an edge in common, but there is no containment relationship between them.
@@
Cycle $C^{\prime}$ is demanding if its red-green coloring obeys Rule 1 and non-demanding otherwise. For an example of application of the coloring rule, see Figure 13(a).
Let $C$ be a cycle incident to the external face $f_{\operatorname{ext}}$, i.e., sharing at least an edge with $f_{\operatorname{ext}}$. The twin of $C$, denoted by $C^{t}$, is the 3-legged cycle having the same legs of $G$. Notice that also $C^{t}$ is a cycle incident to $f_{\operatorname{ext}}$. Let $I(G)$ be the set of demanding cycles of $G$ that are intersecting.
Lemma C.2 ([6]).
@@
See for example Figure 13(b), which depicts a orthogonal drawing with no bends and with minimum dummies of the graph in Figure 13(a), where dummies are represented with a cross. If dummies are removed with bends the drawing has the minimum number of bends.
C.2 A Better Lower Bound for tbn(G)
In this section we refine a lower bound for $tbn(G)$ proving Lemma 4.2. In Section C.3 we prove that such lower bound is tight. Before proving Lemma 4.2 we prove and observe intermediate results.
Observe that, for each unbent collection $\mathcal{C}$ of $G$, $|\mathcal{C}|\geq 2$, as the external face of $G$ is a 0-legged cycle, and by Lemma 4.1, we have to place at least four dummies in $G$. We introduce the following corollary of Lemma C.3.
Corollary C.4.